EVA Report February 1st

[category EVA-report]

Crew 240 EVA Report 02-01-2022

EVA #1

Author: Clément Plagne

Purpose of EVA: First in-sim EVA experience, acquaintance with procedures

Start time: 10:03

End time: 11:00

Narrative: Pre-EVA briefing went smoothly, and we drove to the spot safely and slowly. Finding my bearings with the map and getting my confidence back took a smidge more time than I expected, but we arrived at Marble Ritual without any delay. We walked on Marble Ritual, and a little bit around its rocky formation to get some bearings in the suits. Pierre felt confident enough in the suit to try and take out his drone, which I approved for a quick flight and a picture. A crew member mentioned that they were beginning to feel some back pain, and we therefore decided to call it for the walk and drive back to MDRS, which went well too. We were pleased to see that the batteries for both rovers stayed very high, as they had depleted fairly quickly on training day.

Destination: Marble Ritual

Coordinates (use UTM NAD27 CONUS): 518800 4250700

Participants: Clément Plagne, François Vinet, Pierre Fabre, Julie Levita

Road(s) and routes per MDRS Map: Entrance Road, Cow Dung Road

Mode of travel: Rovers Spirit and Opportunity, foot.

Journalist Report February 1st

Crew 240 Sol Summary Report 02-01-2022

Sol: 1

Summary Title: Putting ourselves out there, somewhat

Author’s name: Clément Plagne

Mission Status: All is well!

Sol Activity Summary: Big day for EVAs, with two occurring in the same day. Morning EVA went on nominally, with crew members getting well acquainted with the procedures and the material at hand. Afternoon EVA had a few issues and required us to get back to the Hab once to fix a headset issue, but we still managed to achieve the objectives set. Meanwhile inside, work continued (with success!) on our 3D printer and for water conservation experiments.

Look Ahead Plan: We have a big EVA tomorrow for setting up our atmospheric experiments, which may take a while if we’re not all well prepared.

Anomalies in work: Crew members have been reporting discomfort with the headsets during EVA, with one microphone working poorly. We’ve put in a request for a new mic, and comfort issues we’ll try to alleviate ourselves. Spacesuit Seven has been spotty and turned off for a few seconds many times over an EVA. François has details of the issue in his Ops report

Weather: Cloudy but (relatively speaking) warm.

Crew Physical Status: Sore from our daily exercise, but no worse for wear

EVA: Two performed (EVA 1: 10:05 to 11:00, EVA 2: 14:41 to 16:25)

Reports to be filed: Operations Report, HSO report (with temperatures), Pictures, Journalist’s report, Sol Summary, EVA Request, EVA Reports (2)

Support Requested: Information about whether or not we can attempt a fix on spacesuit 7 during the coming day, or if Scott thinks it’s a bad idea. Request to use another suit for the next EVA.

We assumed there was a precision scale at the Science Dome but haven’t been able to find it. We read that in a fairly old version of the Science Dome inventory. Is it still there somewhere?

Copyright © The Mars Society. All rights reserved. | Main Site